OpenStudy (anonymous):

Ok i know this is easy but it confuses me when it says w and t vary inversely. LIke i dont know how i should set this up can you explain? Suppose that w and t vary inversely and that t = 1/5 when w = 4. Write a function that models the inverse variation, and find t when w = 9.

OpenStudy (anonymous):

@Disco619 @annas help? @Opcode @beccaboo333

OpenStudy (beccaboo333):

Sorry, I'm of no help here.

OpenStudy (anonymous):

w = 4, as w = 9. w=4 and k= 20, so k/w should be 20/4 = 5.

OpenStudy (anonymous):

I'm 50% sure.

OpenStudy (anonymous):

So when i set it up it would look like this?

OpenStudy (anonymous):

t=k/w?

OpenStudy (anonymous):

Probably.

OpenStudy (anonymous):

when, w = 4 t = 1/5 w = k/t 4 = k/(1/5) 4 = k/5 k = 20. knowing this substitute the new W and the K to the equation solving for Tnew w = k/t t = k/w t = 20/4 get it?

OpenStudy (anonymous):

See! Totally what I said. Just smarter. ♥ Thank you Annas.

OpenStudy (anonymous):

Welcome :D

OpenStudy (anonymous):

Ok i get it now thanks :)

OpenStudy (anonymous):

^_^ my pleasure

OpenStudy (anonymous):

one more medal & I'm on 200 ^_^

OpenStudy (anonymous):

wait! t=20/4 is not an answer choice D:

OpenStudy (anonymous):

my choices are t = one over five w; four over forty-five t = one over five w ; one-fifth t = one over twenty w; one over 80 t = four over five w; four over forty-five

OpenStudy (anonymous):

sorry we need to find when W= 9 so t = 20/9

OpenStudy (anonymous):

oh ok i think i got it now :)

OpenStudy (anonymous):

okies :)

OpenStudy (anonymous):

can I have a cookies now ^_^

OpenStudy (anonymous):

*gives cookies* :)

OpenStudy (anonymous):

yay xD